# Understanding Practical Membership Privacy of Deep Learning

### v1

@@ 6 MIA vulnerability for data samples
This Section extends our investigation from the data set level properties studied in Section 5 to individual data sample level properties. Previously Yu et al. (2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes.
We will focus on a fine-tuned last layer classifier (Head) trained on top of a ViT-B, pre-trained on ImageNet-21k Russakovsky et al. (2015). We investigate two data sample properties that could potentially influence the vulnerability (tpr at fpr) of data points:
@@
  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting Feldman & Zrnic (2021); Koskela et al. (2023). In individual accounting a larger sum of gradient norms results in a higher individual privacy loss. Our hypothesis is that this holds also for MIA vulnerability in non-DP settings.
@@
  • Vector space distances: The pre-trained feature extractors project images into a vector space in which one can compute distances between individual data samples. Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
@@
We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise. For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples. We then compute the tpr at fpr for each bin. We finally report for each data property bin the percentile of the tpr at fpr.
@@
Figure 4: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of data sample properties (binned among a seed). The results have been obtained over 6 seeds with $S=256$ (Patch Camelyon) and $S=64$ (CIFAR10).
Figure 4 shows the observed relationships between the MIA vulnerability and the average individual gradient norms and the vector space distance. We find that:
  • Per-sample gradients: The average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends how many steps are taken into account. Generally, we observe that the correlation is stronger when we only consider the gradients at the end of the training (either the last 20 or 5 steps). This is different from the individual accounting methods that rely on the whole gradient trace for DP accounting.
  • Vector space distances: The distance of data samples to their nearest neighbor correlates with the vulnerability for CIFAR10, but not for Patch Camelyon. We leave a more comprehensive study of this correlation and its limitations for future work.
@@ 7 Discussion
As DP reduces the utility of models, it is important to understand also when it is necessary. Our results on MIA vulnerability as a function of data set properties display a strong power law dependence on the number of examples per class. This leads to an unfortunate dilemma: DP seems most important in the small data regime where it likely reduces the model utility the most.
@@
Due to computational constraints, we only report results on transfer learning with fine-tuning only a part of the model.
@@ Acknowledgements
Appendix A Background
Differential privacy (DP; Dwork et al., 2006b) is the gold standard for formalizing privacy guarantees. $(\epsilon,\delta)$-DP (Dwork et al., 2006a) has a privacy budget consisting of $\epsilon\geq 0$ and $\delta\in[0,1]$, where smaller values correspond to a stronger privacy guarantee. We refer to Dwork & Roth (2014) for a comprehensive intro to DP.
DP deep learning utilizes DP-SGD (Rajkumar & Agarwal, 2012; Song et al., 2013; Abadi et al., 2016), which is the DP adaptation of stochastic gradient descent. DP-SGD clips per-sample gradients to a pre-determined norm to limit the influence of each sample and adds Gaussian noise the sum of the clipped gradients.
Clipping the gradients and adding Gaussian noise in DP-SGD typically comes with a cost in accuracy. This decrease in accuracy can be alleviated through transfer-learning. Publically available pre-trained models are fine-tuned with DP-SGD on sensitive downstream data sets. Both in NLP (Li et al., 2022; Yu et al., 2022) and computer vision (Kurakin et al., 2022; De et al., 2022; Mehta et al., 2022; Cattan et al., 2022; Tobaben et al., 2023) transfer-learning significantly closes the gap between non-private and private training. Tramèr et al. (2022) critically discuss this line of work and the implication of relying on pre-trained models.
Appendix B Training details
B.1 Parameterization
We utilize pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K data set (Russakovsky et al., 2015). We download the feature extractor checkpoints from the respective repositories.
@@ B.1 Parameterization
  • FiLM: In addition to the linear layer from Head, we fine-tune parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network. While a diverse set of adapters has been proposed, we utilize FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
B.2 Hyperparameter tuning
Our hyperparameter tuning is heavily inspired by the comprehensive few-shot experiments by Tobaben et al. (2023). We utilize their hyperparameter tuning protocol as it has been proven to yield SOTA results for (DP) few-shot models.
@@ B.2 Hyperparameter tuning
noise multiplier Based on target $\epsilon$
B.3 Datasets
Table A2 shows the used data sets in the paper. We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner. Additionally, we add CIFAR10 that is not part of the original VTAB benchmark.
@@ B.3 Datasets
Appendix C Additional results
@@
(a) tpr at $\textsc{fpr}=0.1$
(b) tpr at $\textsc{fpr}=0.01$
(c) tpr at $\textsc{fpr}=0.001$
Figure A.1: MIA vulnerability as a function of privacy budget ($\epsilon$ at $\delta=10^{-5}$) when attacking different backbones fine-tuned on CIFAR10 and CIFAR100 at $S\in\{25,50\}$. The solid line displays the median and the error bars the minimum of the lower bounds and the maximum of the upper bounds for the Clopper-Pearson CIs over 15 seeds. The DP bounds are the theoretical DP bounds based on Kairouz et al. (2015) where the naive DP bound is computed with $\delta=10^{-5}$ and the tight DP bound over an interval of deltas and their respective $\epsilon$.
C.1 Additional results for Section 5
@@
C.1.1 Vulnerability as a function of shots
@@
Figure A.2: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream data sets. The errorbars display the min and max Clopper-Pearson CIs over 6 seeds and the solid line the median.
@@
C.1.2 Vulnerability as a function of the number of classes
@@
C.1.3 Predicting data set vulnerability as function of $S$ and $C$
Table A7: Results for fitting Equation 3 with statsmodels Seabold & Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.924 0.898 -0.237 0.008 -30.994 $0$0.252 -0.222 0.01 0.937 0.839 -0.380 0.012 -31.818 $0$0.403 -0.356 0.001 0.917 0.740 -0.492 0.019 -26.143 $0$0.530 -0.455 $\beta_{C}$ (for $C$) 0.1 0.924 0.898 0.095 0.014 6.992 0.000 0.068 0.121 0.01 0.937 0.839 0.226 0.021 10.680 0.000 0.184 0.268 0.001 0.917 0.740 0.357 0.033 10.677 0.000 0.291 0.423 $\beta_{0}$ (intercept) 0.1 0.924 0.898 -0.008 0.029 -0.274 0.785 -0.064 0.049 0.01 0.937 0.839 -0.224 0.045 -5.006 $0$0.312 -0.136 0.001 0.917 0.740 -0.486 0.071 -6.879 $0$0.626 -0.347
@@ C.1.3 Predicting data set vulnerability as function of $S$ and $C$
Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 3 fitted Table A3 (ViT-B). The dots show the median tpr for the train set (ViT-B; Table A3) and the test set (R-50; Table A4) over six seeds. Note that the dots for $C=10$ are for EuroSAT.

### v2

@@ 5 MIA Vulnerability For Data Samples
This section extends our investigation from the data-set-level properties (Section 3) to individual data-sample-level properties. Previously Yu et al. (2023) have studied correlations between classes and privacy risk before, but in comparison to their study, we look at data samples independent from classes. We investigate two properties that could potentially influence the vulnerability of data points:
@@
  • Per-sample gradients: Per-sample gradients form the basis for DP individual accounting (Feldman and Zrnic, 2021; Koskela et al., 2023). Under individual Gaussian DP (GDP, Dong et al., 2022) accounting (Koskela et al., 2023), the cumulative $\mu$-GDP privacy loss of individual $i$ with individual gradient norms $g_{i,t}$ is $\mu_{\text{i,tot}}\propto\sqrt{\sum_{t}g_{i,t}^{2}}$. Our hypothesis is that this similarly holds also for MIA vulnerability in non-DP settings.
@@
  • Feature space distances: The pre-trained feature extractors project images into a feature space in which one can compute distances between individual data samples. Our hypothesis is that data samples that are further away from other data samples are easier to distinguish.
5.1 Experimental Setup
Our model training is done as in Section 3.1. The target models are fine-tuned pre-trained ViT-B (Head) and R-50 (Head) models. Below we describe how we measure the sample properties and individual tpr at fpr.
Recording of gradient norms During the training of the target models we record the per-sample gradient norm for every sample at every step (independent of being part of the current batch) efficiently using Opacus (Yousefpour et al., 2021). The recording of the gradient norms is identical to the DP individual accounting, with the exception that we record the gradient norms of out-samples as well. In our analysis, we are utilising the square root of the sum of squared gradient norm per sample over deciles of steps (i.e. the sum of squared individual gradient norms over the first 10% of steps, second 10% of steps, …, last 10% of steps).
$G_{i}^{(\alpha)}=\sqrt{\sum_{t=\lfloor T(\alpha-0.1)\rfloor}^{\lfloor T\alpha\rfloor}g_{i,t}^{2}},\;\alpha=(0.1,0.2,\ldots,1.0)$ (15)
Computation of distance in feature space All samples are projected into a feature space using the pre-trained model. We measure the Euclidean distance of all examples (in and out) to all in-samples in the feature space and report the distance to the nearest in-sample.
@@ 5.1 Experimental Setup
tpr at fpr for individual samples We study individual samples, but aggregate the results over multiple data samples by binning the samples based on their data sample properties to reduce the noise. For each data sample we compute its property and assign the sample to a bin based on its percentile in relation to other samples. We then compute the tpr at fpr for each bin. We finally report for each data property bin the percentile of the tpr at fpr.
Throughout our experiments we noticed that the amount of data points needs to be sufficiently large, in order to allow for an appropriate estimation of the tpr at fpr. Thus, we use a sufficient amount of shots ($S$) so that the size of in- and out-set combined is at least $10002|\mathcal{D}|=$ 1000 (see Table A10).
5.2 Individual Gradient Norms Are Predictive of Individual MIA Vulnerability
Figure 5 shows the observed relationships between the MIA vulnerability and the average individual gradient norms at $\textsc{fpr}=0.001$. When plotting the data for all deciles of steps at once, we find that the average gradient norms are correlated with the vulnerability in all our experiments, but the strength of the correlation depends on the data set. The $R^{2}$ (coefficient of determination) compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor (corresponding to a scenario where all points are on the diagonal line). See more results in Table A10 and Figure A.6 in the Appendix.
Figure 5: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed). Data obtained over six seeds. The $R^{2}$ score compares the observed relationship to a scenario where individual gradient norms would be a perfect predictor.
Which parts of the training are more predictive?
In Figure 6, we examine whether the gradient norms at different stages of training are a better predictor of vulnerability by computing the ratio between the $R^{2}$ per iteration decile and $R^{2}$ over the whole training. For most data sets the ratio is constantly at a value of roughly 1, which suggests that gradient norms over even short segments can be as predictive as the norms of the whole trace.
Furthermore, there are deviations that indicate that for some data sets, gradient norms during the early iterations can be relatively more predictive than the norms of the entire trace. This is interesting, as the DP bound would directly depend on the norms of the entire trace.
Figure 6: Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training. A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predicitive.
5.3 Distance to Neighbors in Feature Space of Pre-trained Model
Figure 7 shows the observed relationships between the MIA vulnerability and distance to the nearest neighbor in the feature space. The distance of data samples to their nearest neighbor correlates with the vulnerability all analysed data sets but Patch Camelyon and Resics45. The observation holds over all fpr and for R-50 (Head) as well, refer to Figures A.9 and A.10 in the Appendix.
@@ 5.3 Distance to Neighbors in Feature Space of Pre-trained Model
Figure 7: MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
@@ 6 Discussion
Under the GDPR and similar legal regimes, machine learning (ML) models trained on personal data that memorise the data are personal data and need to be carefully protected. Our work analyses in which cases trained models would most likely be personal data and in which cases they might not be. This will help in evaluating the risk of different kinds of models, favouring less risky models when possible and paying extra attention to more risky cases.
As the best means of protecting privacy, DP, reduces the utility of models, it is important to understand when it is necessary. Our results highlight that models are most vulnerable to MIA when the number of examples per class is low. This leads to an unfortunate dilemma: DP seems most important in the regime where it likely reduces the model utility the most.
Our model singles out the number of examples per class $S$ as the key predictor of model vulnerability. While our results are based on balanced data sets, it seems likely that minority classes and outliers would be more vulnerable as predicted by our model given their smaller $S$. Studying the vulnerability in unbalanced data and in the presence of outliers is an important area of future work.
@@
Our work measures the vulnerability through running one particular MIA. There is no guarantee that models that are not vulnerable to this attack might not be vulnerable to some stronger attack. Formal bounds on attack susceptibility would require something like DP.
Due to computational constraints, our predictive model is based only on transfer learning with fine-tuning only a part of the model. Limited comparison of our model predictions to previous results indicates it may underestimate their vulnerability.
Broader Impact
Our work systematically studies factors influencing privacy risk of trained ML models. This has significant positive impact on users training ML models on personal data by allowing them to understand and limit the risks.
Reproducibility
We provide the code for reproducing the experiments in the supplementary material and will publish the code in an open repository after acceptance of the paper. All used pre-trained models and data sets are publicly available.
@@ Acknowledgements
Appendix A Training details
A.1 Parameterization
We utilise pre-trained feature extractors BiT-M-R50x1 (R-50) (Kolesnikov et al., 2020) with 23.5M parameters and Vision Transformer ViT-Base-16 (ViT-B) (Dosovitskiy et al., 2021) with 85.8M parameters, both pretrained on the ImageNet-21K data set (Russakovsky et al., 2015). We download the feature extractor checkpoints from the respective repositories.
@@ A.1 Parameterization
  • FiLM: In addition to the linear layer from Head, we fine-tune parameter-efficient FiLM (Perez et al., 2018) adapters scattered throughout the network. While a diverse set of adapters has been proposed, we utilise FiLM as it has been shown to be competitive in prior work (Shysheya et al., 2023; Tobaben et al., 2023).
A.1.1 Licenses and Access
The licenses and means to access the model checkpoints can be found below.
A.2 Hyperparameter tuning
@@
A.3 Data sets
Table A2 shows the data sets used in the paper. We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner. Additionally, we add CIFAR10 which is not part of the original VTAB benchmark.
@@ A.3 Data sets
A.3.1 Licenses and Access
The licenses and means to access the data sets can be found below. We downloaded all data sets from TensorFlow datasets https://www.tensorflow.org/datasets but Resics45 which required manual download.
A.4 Compute Resources
All experiments but the R-50 (FiLM) experiments are run on CPU with 8 cores and 16 GB of host memory. The training time depends on the model (ViT is cheaper than R-50), number of shots $S$ and the number of classes $C$ but ranges for the training of one model from some minutes to an hour. This assumes that the images are passed once through the pre-trained backbone and then cached as feature vectors. The provided code implements this optimization.
In Section 5 we record the per-sample gradients during the training. This makes the training much more expensive despite the use of efficient per-sample gradient computation using Opacus Yousefpour et al. (2021). The training of the considered models does still not exceed one hour per model as we only consider settings with relatively low $S$ for this part of the experiments.
The R-50 (FiLM) experiments are significantly more expensive and utilise a NVIDIA V100 with 40 GB VRAM, 10 CPU cores and 64 GB of host memory. The training of 257 shadow models then does not exceed 24h for the settings that we consider.
We estimate that in total we spend around 7 days of V100 and some dozens of weeks of CPU core time but more exact measurements are hard to make.
Appendix B Additional results
@@
B.1 Additional results for Section 3
@@
B.1.1 Vulnerability as a function of shots
@@
Figure A.1: MIA vulnerability as a function of shots (examples per class) when attacking a pre-trained ViT-B and R-50 Head trained without DP on different downstream data sets. The errorbars display the minimum and maximum Clopper-Pearson CIs over six seeds and the solid line the median.
@@
B.1.2 Vulnerability as a function of the number of classes
@@
B.1.3 Data for FiLM and from scratch training
Table A7: MIA vulnerability data used in Figure 3(b). Note that the data from Carlini et al. (2022) is only partially tabular, thus we estimated the tpr at fpr from the plots in the Appendix of their paper.
model dataset classes shots source tpr@ tpr@ tpr@ ($C$) ($S$) fpr=0.1 fpr=0.01 fpr=0.001 R-50 FiLM CIFAR10 10 50 This work 0.482 0.275 0.165 (Krizhevsky, 2009) CIFAR100 100 10 Tobaben et al. (2023) 0.933 0.788 0.525 (Krizhevsky, 2009) 25 Tobaben et al. (2023) 0.766 0.576 0.449 50 Tobaben et al. (2023) 0.586 0.388 0.227 100 Tobaben et al. (2023) 0.448 0.202 0.077 EuroSAT 10 8 This work 0.791 0.388 0.144 (Helber et al., 2019) Patch Camelyon 2 256 This work 0.379 0.164 0.076 (Veeling et al., 2018) Pets 37 8 This work 0.956 0.665 0.378 (Parkhi et al., 2012) Resics45 45 32 This work 0.632 0.379 0.217 (Cheng et al., 2017) from scratch CIFAR10 10 2500 Carlini et al. (2022) 0.300 0.110 0.084 (Krizhevsky, 2009) (wide ResNet) CIFAR100 100 250 Carlini et al. (2022) 0.700 0.400 0.276 (Krizhevsky, 2009)
B.1.4 Predicting data set vulnerability as function of $S$ and $C$
This section provides additional results for the model based on Equation 2
Table A8: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 0.0001 0.852 0.618 -0.741 0.035 -21.467 $0$0.809 -0.673 0.00001 0.837 0.404 -0.836 0.045 -18.690 $0$0.924 -0.748 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 0.0001 0.852 0.618 0.363 0.065 5.616 0.000 0.236 0.491 0.00001 0.837 0.404 0.569 0.085 6.655 0.000 0.400 0.737 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022 0.0001 0.852 0.618 -0.303 0.138 -2.202 0.029 -0.575 -0.032 0.00001 0.837 0.404 -0.615 0.180 -3.414 0.001 -0.970 -0.260
Figure A.3 shows the performance for all considered fpr.
[…]
Figure A.3: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation 2 fitted Table A3 (ViT-B). The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds. Note that the triangles and dots for $C=10$ are for EuroSAT.
B.2 Simpler variant of the prediction model
[…]
The general form this variant can be found in Equation A1, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
$\log_{10}(\textsc{tpr})=\beta_{S}\log_{10}(S)+\beta_{C}\log_{10}(C)+\beta_{0}$ (A1)
Table A9 provides tabular results on the performance of the variant.
Table A9: Results for fitting Equation A1 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001,0.0001,0.00001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 0.0001 0.865 0.628 -0.697 0.031 -22.274 $0$0.758 -0.635 0.00001 0.862 0.400 -0.802 0.040 -20.311 $0$0.880 -0.725 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 0.0001 0.865 0.628 0.371 0.059 6.328 0.000 0.255 0.486 0.00001 0.862 0.400 0.580 0.076 7.679 0.000 0.431 0.729 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161 0.0001 0.865 0.628 -0.387 0.125 -3.104 0.002 -0.633 -0.141 0.00001 0.862 0.400 -0.683 0.159 -4.288 $0$0.996 -0.369
Figure A.4 plots the performance of the variant similar to Figure 3(a) in the main text.
@@ B.2 Simpler variant of the prediction model
Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation A1 fitted Table A3 (ViT-B). The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds. Note that the triangles and dots for $C=10$ are for EuroSAT.
B.3 Additional results for Section 4
This section provides additional results for Section 4.
B.4 Rates of In- and Out-distribution Parameters as Function of data set Properties
LiRA assumes the attacker can build Gaussian distributions of losses for each data point $x$ through the training of shadow models (see Section 2). We denote the distributions as following:
  • In-distribution: Distribution of the losses of shadow models using $x$ for training ($x\in\mathcal{D}$)
  • Out-distribution: Distribution of the losses of shadow models not using $x$ for training ($x\notin\mathcal{D}$)
In Figure A.5 we visualise the changes of the two distributions as functions of $S$ and $C$.
(a) Distribution means
(b) Distribution variances
(c) Distribution standard deviations
(d) Difference of Means
(e) Difference of Means / std
(f) Difference of Means / var
Figure A.5: Means of observed statistics of the in- and out-distributions for the simplified model with in-class $\sigma=0.1$. The statistics are obtained for $C=2$.
B.5 Additional results for Section 5
This section provides additional results for Section 5. Figures A.6 and A.7 provide additional data for all considered fpr and for both ViT-B (Head) and R-50 (Head). Table A10 the $R^{2}$ scores for ViT-B (Head) and the experiment settings for ViT-B (Head) and R-50 (Head). Figure A.8 the relative $R^{2}$ scores for R-50 (Head).
B.5.1 MIA vulnerability as a function of individual gradient norms
Figure A.6: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed). The results have been obtained over six seeds. The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
Figure A.7: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of individual gradient norms (binned among a seed). The results have been obtained over six seeds. The number in the bottom right is the $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
Table A10: ViT-B (Head): $R^{2}$ (coefficient of determination) score when comparing the observed relationship between tpr at fpr and individual gradient norms to a scenario where individual gradient norms are a perfect predictor of individual MIA vulnerability (corresponding to a scenario where all points are on the diagonal line).
dataset shots ($S$) classes ($C$) data set size (|$\mathcal{D}$|) $R^{2}$ at fpr=0.1 $R^{2}$ at fpr=0.01 $R^{2}$ at fpr=0.001 Patch Camelyon (Veeling et al., 2018) 256 2 512 0.98 0.97 0.81 CIFAR10 (Krizhevsky, 2009) 64 10 640 0.98 0.98 0.92 EuroSAT (Helber et al., 2019) 64 10 640 0.98 0.97 0.92 Pets (Parkhi et al., 2012) 32 37 1184 0.95 0.93 0.89 Resics45 (Cheng et al., 2017) 32 45 1440 0.95 0.97 0.97 CIFAR100 (Krizhevsky, 2009) 16 100 1600 0.89 0.89 0.86
Figure A.8: R-50 (Head): Ratio between $R^{2}$ per iteration decile and $R^{2}$ over the whole training. A ratio $>1$ means that the decile is as more predictive than whole trace and a ratio $<1$ means that it is less predictive.
B.5.2 MIA vulnerability as a function of feature space distance
This section provides additional data regarding the individual MIA prediction using feature space distances.
Figure A.9: ViT-B (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
Figure A.10: R-50 (Head): MIA vulnerability (tpr at fpr; percentiles among a seed) as a function of feature space distance (binned among a seed). The results have been obtained over six seeds.
Appendix C Missing proofs of Section 4.2
See 1
Proof.
Recall
fpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{fpr}}}>\tau\mid x\notin\mathcal{D}_{\textrm{target}}\right)$ (A2)
tpr $=\Pr\left(\underbrace{-\frac{1}{2\sigma^{2}}(\mu_{\textrm{in}}-\mu_{\textrm{out}})\left[-2t_{x}+(\mu_{\textrm{in}}+\mu_{\textrm{out}})\right]}_{\lambda_{\textsc{tpr}}}>\tau\mid x\in\mathcal{D}_{\textrm{target}}\right).$ (A3)
We denote the class means for the $m$th in/out shadow data sets with $r_{y,m}^{(\textrm{in})}$ and $r_{y,m}^{(\textrm{out})}$
$r_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\sum_{i=1}^{S-1}x_{i,m,y}+x\right),\;r_{y,m}^{(\textrm{out})}=\frac{1}{S}\left(\sum_{i=1}^{S}x_{i,m,y}\right),$ (A4)
where $x_{i,m,y}$ denotes the $i$th element of the $m$th shadow data set with label $y$. The corresponding scores $s_{y,m}^{(\textrm{in})}=\langle x,r_{y,m}^{(\textrm{in})}\rangle$ and $s_{y,m}^{(\textrm{out})}=\langle x,r_{y,m}^{(\textrm{out})}\rangle$ which can be written as
$s_{y,m}^{(\textrm{in})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S-1}x_{i,m,y}\rangle+\langle x,x\rangle\right)=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle+\langle x,x-x_{S,m,y}\rangle\right)$ (A5)
$s_{y,m}^{(\textrm{out})}$ $=\frac{1}{S}\left(\langle x,\sum_{i=1}^{S}x_{i,m,y}\rangle\right)$ (A6)
Now, assuming we have arbitrarily many shadow models, we can write the LiRA statistics as
$\mu_{\textrm{in}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{in})}],\;\mu_{\textrm{out}}=\operatorname{\mathbb{E}}_{D_{m}}[s_{y,m}^{(\textrm{out})}]=\langle x,m_{y}\rangle$ (A7)
$\Rightarrow$ $\mu_{\textrm{in}}-\mu_{\textrm{out}}=\frac{1}{S}\langle x,x-\mathbb{E}_{D_{m}}[x_{S,m}]\rangle=\frac{1}{S}\langle x,x-m_{y}\rangle,$ (A8)
where $m_{y}$ is the true center of the $y$th cluster. For $\sigma$ we have
$\sigma^{2}$ $=\operatorname{Var}(s_{y,m}^{(\textrm{out})})$ (A9)
$=\frac{1}{S^{2}}\sum_{i=1}^{S}\operatorname{Var}(\langle x,x_{i,m}\rangle)$ (A10)
$=\frac{1}{S}\operatorname{Var}(\langle x,x_{m}^{(j)}\rangle)$ (A11)
$=\frac{1}{S}\operatorname{Var}(x_{m}^{(j)})||x||^{2}=\frac{1}{S}s^{2}||x||^{2},$ (A12)
where $x_{m}^{(j)}$ denotes the $j$th dimension of the vector $x_{m}$ and $s$ is the std of each dimension of $x_{m}$. Now we can write the $\lambda_{\textsc{fpr}}$ as
$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(-2t_{x}+2\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A13)
In the fpr case, we have that $x\notin\mathcal{D}_{\textrm{target}}$, therefore for the $t_{x}$ we have
$t_{x}=\langle x,r_{y}^{\textrm{(target)}}\rangle\sim\langle x,m_{y}\rangle+\frac{s}{\sqrt{S}}\langle x,z\rangle,$ (A14)
where $z\sim N(0,I_{d})$. Substituting this into $\lambda_{\textsc{fpr}}$ we have
$\lambda_{\textsc{fpr}}=-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A15)
We see that $\lambda_{\textsc{fpr}}$ evolves as $O(1/\sqrt{S})$.
Next, we will focus on the $\lambda_{\textsc{tpr}}$. The expression remains otherwise the same as for $\lambda_{\textsc{fpr}}$, but since $x\in\mathcal{D}_{\textrm{target}}$ the $t_{x}$ changes to
$t_{x}\sim\langle x,m_{y}\rangle+\frac{1}{S}\langle x,x\rangle+s\frac{1}{\sqrt{S}}\sqrt{1-\frac{1}{S}}\langle x,z\rangle.$ (A16)
Hence, for large enough $S$ we have
$\lambda_{\textsc{tpr}}\approx-\frac{1}{2s^{2}||x||^{2}}\langle x,x-m_{y}\rangle\left(\frac{2s}{\sqrt{S}}\langle x,z\rangle-\frac{2}{S}||x||^{2}+\frac{1}{S}\langle x,x-m_{y}\rangle\right).$ (A17)